OpenStudy (anonymous):

What is the solution of the equation 4x + 1 = 21? Round your answer to the nearest ten-thousandth.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

4x+1=21 -1. -1 4x=20 /4. /4 X=20/4 X=5